Ottawa-INVEST is offering an investment that pays an APR of 5% with continuous compounding. What is the effective annual rate of return provided by this investment?

Question 23 options:

5.50%

4.12%

5.13%

5.00%

Not enough information.
Answer #1

APR = 5% with continuous compounding

EAR = e^APR -1

Effective Annual Rate of return = e^APR -1 = e ^ 5% -1 = 5.13%

Answer is 5.13%